Dyrdek Enterprises has equity with a market value of $10.9 million and the market value of debt is $3.60 million. The company is evaluating a new project that has more risk than the firm. As a result, the company will apply a risk adjustment factor of 1.7 percent. The new project will cost $2.22 million today and provide annual cash flows of $581,000 for the next 6 years. The company's cost of equity is 11.11 percent and the pretax cost of debt is 4.89 percent. The tax rate is 35 percent. What is the project's NPV?

Multiple Choice

  • $219,241

  • $230,173

  • $249,256

  • $375,414

  • $512,072
Answer #1

Answer :

Market value of equity = 10,900,000

Market value of debt = 3,600,000

Total market value = 10,900,000 + 3,600,000 = 14,500,000

Weight of equity = 10,900,000 / 14,500,000 = 0.752

Weight of debt = 3,600,000 / 14,500,000 = 0.2483

Pre tax cost of debt = 4.89%

After tax cost of debt = 0.0489 ( 1 - 0.35 )

After tax cost of debt = 0.031785 (or) 3.1785%

WACC = 0.752 * 0.1111 + 0.2483 * 0.031785

= 0.0835472 + 0.0078922

WACC = 0.0914394 (or) 9.144%.

Since the project is riskier , WACC of project = 9.144 + 1.7 = 10.844%

NPV = Present value of cash inflows - Present value of cash outflows

NPV = -2,220,000 + 581,000 / ( 1+0.10844 )1 + 581,000 / ( 1+0.10844 )2 + 581,000 / ( 1+0.10844 )3 + 581,000 / ( 1+0.10844 )4 + 581,000 / ( 1+0.10844 )5 + 581,000 / ( 1+0.10844 )6

NPV = $249,256.